OpenStudy (anonymous):

Can someone help me please refresh my memory how to solve this problem? Thanks in advance! 5x/4+2=1/3

OpenStudy (anonymous):

Ok Im sorry they give the answers -4/3 -23/15 19/15 4/5 Can you add 4 which is a denominator to 2 which is a whole number? Thanks for helping me so far:)

OpenStudy (anonymous):

I'm confused on how you laid that out...sorry I am still learning this.

OpenStudy (anonymous):

5x/4+2=1/3 So subtract 2 over to be left with 5x/4 on the left side. 5x/4=-5/3 Then multiply 4 to get rid of the denominator. You're left with 5x on the left side which is equal to -20/3. 5x=-20/3 Divide by 5 to get x alone. (-20/3) divide by 5 is equal to -4/3 x=-4/3

OpenStudy (anonymous):

Thank you so much!:))